OpenStudy (anonymous):

30 more than 10 divided by the sum of 2 and 3

OpenStudy (anonymous):

30+10/(2+3) or (30+10)/(2+3) , depends on how it's asked.

OpenStudy (anonymous):

they asking to write an expression for the situation

OpenStudy (anonymous):

It's not an inequality...

OpenStudy (anonymous):

hmm ok, will delete

OpenStudy (anonymous):

write an expression for the situation 30 more than 10 divided by the sum of 2 and 3

OpenStudy (anonymous):

30 + 10/(2+3) = 30 + 10/5 = 32

OpenStudy (anonymous):

simplify the expression

OpenStudy (anonymous):

yes but they asking divided

OpenStudy (anonymous):

order of operations? 30 more than 10 divided by the sum of 2 and 3 30 more than (10 divided by the sum of 2 and 3) 30 more than (10 divided by the (sum of 2 and 3) ) 30 more than (10 divided by 5) 30 more than 2 32

OpenStudy (anonymous):

good choice :) Kidding aside, it is not worded well... hate these vague questions!!1

OpenStudy (anonymous):

(same, less the 1 on the end)

OpenStudy (anonymous):

two good choices... you're on a roll!

OpenStudy (anonymous):

@mygirls56 does this make any sense? It's just a matter of working it out step by step

OpenStudy (anonymous):

yes JakeV8 THANK YOU

OpenStudy (anonymous):

your welcome... good ideas from all above, but glad you could work through it.

OpenStudy (anonymous):

YES SO TRUE